OpenStudy (anonymous):

a basketball player signs a contract with a starting salary of $850,000 per year and an annual increase of 4.5% beginning the second year. What will the athlete's salary be in the sixth year?

OpenStudy (anonymous):

the point of these problems is to get you to understand how to do it on your graphing calculator.

OpenStudy (anonymous):

Irawrd is right, but can it can be done algebraically. Percent is out of 100 so 4.5%= .045 You could multiply 850000 x .045 = 38250+850000= 888250 x .045 =39971.25 + 888250 = 928221.25 and on and on to six times. That's the hard way. The easy way is to just put it in the graphing calculator.